OpenStudy (anonymous):

Write the equation of the line whose slope is -3/7 and y-intercept is -4; write the answer in slope-intercept form

OpenStudy (anonymous):

y= -3/7x -4

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

y=mx+c where m is the slope and c is the y-intercept y=-3/7(x)+(-4) y=-3x/7 - 4 y=-3x-28/7 7y = -3x - 28 3x+7y + 28 =0